What is the inverse function of f(x)=x+1?
Answer by stanbon(75887)

inverse function of f(x)=x+1
-------------
Rewrite as y=x+1.
Interchange x and y to get:
x=y+1
Solve for y to get:
y=x-1
That is the inverse.
Cheers,
Stan H.